Calculate the molar mass of the following substances.
      
        (d) Hydrochloric acid, HCl
 

Grade:12th pass

1 Answers

Harshit Singh
2 years ago
Dear Student


Hydrochloric acid, HCl = 1 × 1 + 1 × 35.5
= 1 + 35.5 = 36.5 g


Thanks